Please provide explanation to the correct answer.

Question #12

Selected accounts with some debits and credits omitted are presented as follows:

                                                      Work in Process

Oct. 1

Balance

20,000  

Oct. 31

Goods finished

X

      31

Direct materials

96,700  

      31

Direct labor

201,000  

      31

Factory overhead

X  

                                                   Finished Goods

Oct. 1

Balance

52,000  

      31

Goods finished

360,000  


If the balance of Work in Process at October 31 is $21,000, what was the amount of factory overhead applied in October?

  $11,300

  $42,300---Incorrect

  $21,300

  $63,300

Question #13

If the amount of actual factory overhead cost incurred exceeds the amount applied, the factory overhead is:

over-applied----Incorrect

  max-applied

  straight-line applied

  under-applied

Question #15

The recording of the jobs shipped and customers billed would include a debit to:

  Cash

  Accounts Payable

  Finished Goods---Incorrect

  Cost of Goods Sold

Question # 24

In process costing, the allocation of current period cost to ending WIP will be based on

  Started and Completed Units EUP

  Ending WIP EUP

  Finished Goods EUP

  Beginning WIP EUP------Incorrect
Answer #1
12)_
Goods finished is $360,000 which is given in the finished goods account.
Goods finished + work in progress at October 31= October 1 balance + Direct materials + Direct labor+ Factory overhead applied
360,000 + 21,000 = 20,000 + 96,700 + 201,000 + Factory overhead applied
Factory overhead applied = 360,000 + 21,000 -20,000-96,700 - 201,000
Factory overhead applied = $63,300
Hence, the correct option is $63,300
13)
Overapplication of overhead - Overhead applied is more than the actual overhead
Underapplication of overhead - Overhead applied is less than the actual overhead.
In the given question overhead applied is less than the actual overhead which shows
that the overhead is underapplied.,
Hence, the correct option is Under-applied.
15)
Entry
Account Titles and Explanation Debit ($) Credit ($)
Cost of good sold xx
Finished goods xx
The above entry includes a debit to cost of goods sold.
Hence, the correct option is cost of goods sold.
24)
Equivalent unit of production of ending work-in-process are computed
with respect to its % of completion and cost during the current period
are allocated based on the ending WIP EUP.
Hence, the correct option is Ending WIP EUP.